OpenStudy (anonymous):

For f(x)= 5 x + 1, find f(-4). A -19 B 1 C -21 D 21 Can you also explain this? I don't quite get it. Thanks :)

OpenStudy (anonymous):

It's A for anyone who was wondering. I still don't understand this... :L

OpenStudy (kirbykirby):

You plug in the value x=-4 into the formula: \(5(-4)+1\) \(=-20+1\) \(=-19\) Since 5x represents a multiplication, you perform a multiplication when you plug in x =-4

OpenStudy (anonymous):

Thank you :)